In one experiment, 1.25g of an ester was completely burnt in an excess of oxygen. 2.84g of carbon dioxide and 1.16g of water were the only products formed. Assuming that the ester contains only the elements carbon, hydrogen and oxygen, calculate the masses of carbon, hydrogen and oxygen in 1.25g of the compound
1
Expert's answer
2020-08-13T07:34:43-0400

Weight of given ester= 1.25 g

Carbon dioxide formed(CO2)= 2.84 g

Water formed (H2O)= 1.16 g

Number of mole of CO2 is = 2.84/44

=.065

That means no.of mole of Carbon is = 0.065

Mass of carbon is = 0.065×12=.78g

Mole of H2O formed= 1.16/18

=.065

No.of mole of Hydrogen is = .065

Mass of hydrogen (H)= .065×1= .065g

Total Mass of ester is=1.25g

Mass of carbon in ester is =.78g

Mass of hydrogen in ester is = .065g

Mass of oxygen in ester is

= 1.25 - 0.845 =.405 g